# Graph of Attacks with Pruning: Optimizing Stealthy Jailbreak Prompt Generation for Enhanced LLM Content Moderation

### v2

@@ Abstract
As large language models (LLMs) become increasingly prevalent, ensuring their robustness against adversarial misuse is crucial. This paper introduces the GAP (Graph of Attacks with Pruning) framework, an advanced approach for generating stealthy jailbreak prompts to evaluate and enhance LLM safeguards. GAP addresses limitations in existing tree-based LLM jailbreak methods by implementing an interconnected graph structure that enables knowledge sharing across attack paths. Our experimental evaluation demonstrates GAP’s superiority over existing techniques, achieving a 20.8% increase in attack success rates while reducing query costs by 62.7%. GAP consistently outperforms state-of-the-art methods for attacking both open and closed LLMs, with attack success rates of $\geq$96%. Additionally, we present specialized variants like GAP-Auto for automated seed generation and GAP-VLM for multimodal attacks. GAP-generated prompts prove highly effective in improving content moderation systems, increasing true positive detection rates by 108.5% and accuracy by 183.6% when used for fine-tuning.  [^1] Warning: This paper contains examples of adversarial prompts that may be offensive to readers. Our implementation is available at https://github.com/dsbuddy/GAP-LLM-Safety.
[…]
@@ 1 Introduction
With the increasing adoption of large-language models (LLMs) across diverse applications, ensuring their reliability and robustness against adversarial misuse has become a critical priority Chao et al. (2023). Jailbreaking techniques, which involve crafting adversarial prompts to bypass an LLM’s safeguards, pose a persistent challenge to AI security and responsible deployment Shen et al. (2024); Mangaokar et al. (2024); Wei et al. (2024); Li et al. (2023); Guo et al. (2024). These methods can induce models to generate harmful, biased, or unauthorized content while avoiding detection by automated moderation systems Perez et al. (2022), highlighting the need for comprehensive diagnostic frameworks to assess and improve foundation model reliability.
@@
GAP-Enhanced Prompt Guard 68.0% 100.0% 100.0% 66.0% 70.0%
Table 1: True positive rate (TPR) comparison of various guardrails detecting prompts generated from multiple jailbreak methods (on AdvBench seeds). Lower TPR indicates better evasion and significant reliability concerns. Jailbreaking prompts generated by TAP and GAP reveal the most critical vulnerabilities across most guardrails. The last row shows how GAP-generated data can be used to enhanced content moderation systems, demonstrating substantially improved detection capabilities against all methods, including GAP itself. Highest TPR values are bolded.
@@
Existing jailbreaking methods fall into three broad categories: (a) white-box attacks, which leverage direct model access for adversarial optimization Zou et al. (2023); Geisler et al. (2024); (b) gray-box attacks, which involve techniques such as backdoor injection or poisoned retrieval Ding et al. (2023); Shi et al. (2023); Zou et al. (2024); Wang and Shu (2023); and (c) black-box attacks, which require only API access and thus represent the most realistic scenario for evaluating model robustness in real-world deployments Wei et al. (2024); Li et al. (2023); Yu et al. (2023); Yuan et al. (2023). Recent advances include AutoDAN-Turbo Liu et al. (2024a), which employs a lifelong learning approach to automatically discover and evolve jailbreak strategies through multi-agent frameworks and strategy libraries. However, AutoDAN-Turbo focuses on long-term strategy accumulation and requires extensive warm-up phases, making it unsuitable as a direct baseline for our work, which addresses immediate structural limitations in prompt exploration efficiency. Notably, the Tree of Attacks with Pruning (TAP) approach Mehrotra et al. (2023) introduced a tree-structured exploration process for iterative prompt refinement, generating increasingly effective adversarial inputs that appear human-like and stealthy. As shown in Table 1, TAP-generated jailbreak prompts consistently demonstrate low detection true positive rate (TPR) when run against recent guardrails, indicating significant vulnerabilities in these safeguard systems that require systematic assessment and improvement.
While TAP demonstrated effectiveness in generating stealthy jailbreaks, we identified several limitations when applying it to thoroughly evaluate model reliability. Primarily, TAP restricts the exploration of prompt refinement to individual paths, with no crossover or shared context across different branches. This isolated approach results in redundant queries and inefficient coverage of the search space for prompt refinement. Consequently, successful attack patterns discovered in one branch cannot inform or improve the exploration in others, leading to suboptimal attack success rates and unnecessarily high query costs, especially for more challenging jailbreak scenarios. To address these limitations in vulnerability assessment, we developed the GAP (Graph of Attacks with Pruning) framework, which: (1) converts the tree-based prompt exploration process into an interconnected graph structure, (2) implements global context maintenance to aggregate successful jailbreak generation strategies, and (3) facilitates graph-based knowledge sharing for more informed prompt refinement.  [^2] Our threat model focuses on forcing LLMs to produce harmful responses through black-box user prompt access only, to account for various LLMs and scenarios where system prompts are inaccessible. As shown in Table 1, GAP achieves significantly higher success rates in bypassing various guardrails while demonstrating superior stealth (lower TPR) compared to TAP. Notably, GAP demonstrates superior evasion capabilities against the Prompt Guard, with a TPR of 16.0% compared to TAP’s 22.0%.
@@
  • Most significantly, we demonstrate how GAP-generated insights can directly improve foundation model reliability through data augmentation and fine-tuning of safeguards. Our experiments demonstrate that GAP-Enhanced Prompt Guard significantly improves detection capabilities across all jailbreak methods, including those identified by GAP itself. As shown in Table 1, the GAP-Enhanced Prompt Guard achieves a TPR of 70.0% against GAP, versus the original Prompt Guard’s 16.0%, demonstrating a substantial improvement in content moderation effectiveness.
@@ 2 Methodology
Algorithm 1 GAP (Graph of Attacks with Pruning)
1:Query $Q$, branching-factor $b$, maximum width $w$, maximum depth $d$
2:Jailbreak prompt $p$ or failure
4:while depth of $G\leq d$ $\triangleright$ Step 3: Iteration
5:     for each leaf node $\ell$ in $G$ do
6:         $C\leftarrow\{\}$triangleright Initialize empty set for conversation histories
7:         for each path from root to a leaf in $G$ do
8:              $h\leftarrow$ Concatenate all $[p,r,s]$ tuples in the path
9:              $C\leftarrow C\cup\{h\}$triangleright Add path history to set
10:         $global\_context\leftarrow$ SortByMaxScore($C$$\triangleright$ Step 1: Build global context
11:         for $j\leftarrow 1$ to $b$ $\triangleright$ Step 1: Child-generation
12:              $p_{j}\leftarrow\mathcal{A}(Q,global\_context)$triangleright Generate prompt using Attacker
13:              $s_{j}\leftarrow$ Retrieve effectiveness of $p_{j}$ based on $global\_context$
14:         $p_{best}\leftarrow\operatorname*{arg\,max}_{j}s_{j}$
[…]
16:         Add child of $\ell$ with prompt $p_{best}$ and history $new\_history$
19:     if successful jailbreak found then return jailbreak prompt
21:return failure
@@ 2.1 GAP (Graph of Attacks with Pruning)
In summary, the GAP core algorithm includes three stages:
@@
  • (Step 2) The pruning step where the judge LLM evaluates branches, removes unsuccessful ones, and focuses effort on variants most effective at eliciting undesired responses (lines 15 and 18).
@@
For the second step, GAP implements a two-phase pruning strategy:
  1. 1.
    Phase 1 (Off-topic pruning): The judge LLM removes branches irrelevant to the original harmful request (line 15).
  2. 2.
    Phase 2 (Highest-scoring pruning): After evaluating target LLM responses, only branches with the highest scores $s_{i,j}=\mathcal{J}(p_{i,j},r_{i,j})$ (up to width $w$) advance to the next iteration (line 18).
For the first step, GAP’s key innovation is its global context $C=\{h_{1},h_{2},\dots,h_{n}\}$ that aggregates successful attack patterns from prior generations across all branches and sequential seeds (lines 4-8). For each prompt node $p$, GAP maintains a history $h_{p}$ of [prompt, response, score] tuples along its refinement path. Unlike TAP’s isolated tree structure, where each seed generates an independent attack path, GAP maintains a unified attack graph where successful strategies are shared and reused. This enables each new seed to leverage patterns observed in previous seeds, resulting in progressively smaller, more efficient attack trees with each sequential seed, as illustrated in Figure 1.
@@
2.1.1 Knowledge Transfer Implementation
GAP’s exploration of prompt generation follows an interconnected graph-structured thought process. The proposed global context enables knowledge transfer through two key mechanisms designed in Step 1 of GAP:
  1. 1.
  2. 2.
The attacker LLM uses this global context when creating jailbreak attempts with two goals: (1) crafting natural-sounding prompts likely to elicit target responses and (2) incorporating effective patterns observed across successful examples in the global context. This guidance to leverage successful patterns enables the attacker to reuse and adapt proven strategies to the current context, improving jailbreak efficiency.
While TAP Mehrotra et al. (2023) represents the closest related work in current literature, it fundamentally differs from our approach by restricting exploration to isolated tree structures. In contrast, GAP’s interconnected graph architecture enables cross-branch knowledge sharing and pattern reuse, as visualized in Figure 1. This structural difference explains GAP’s superior performance in both efficiency and effectiveness, which we quantitatively demonstrate through comprehensive empirical evaluation in Section 3.
2.2 GAP Variants for Different Scenarios
Table 2: Comparison of TAP and GAP variants. While all GAP variants use a graph structure with shared knowledge, they differ in their specific capabilities and the underlying attacker models we choose to use for generating jailbreak prompts.
@@ 2.2.1 GAP-Auto: Auto Seed Generation
While GAP effectively generates sophisticated jailbreak prompts, it initially requires manually crafted seed examples. To eliminate this dependency, we developed GAP-Auto, which automatically generates diverse seed prompts through a two-phase strategy:
@@
Algorithm 2 GAP-Auto Seed Generation
1:High-level content policies
2:$B\leftarrow$ DecomposeIntoBehaviors(content policies)
3:$S_{benign},S_{harmful}\leftarrow\{\},\{\}$
4:for each behavior $b$ in $B$ do
5:     $s_{benign}\leftarrow$ GenerateBenignPrompt($b$)
6:     $s_{harmful}\leftarrow$ GenerateHarmfulPrompt($b$)
7:     $S_{benign}\leftarrow S_{benign}\cup\{s_{benign}\}$
8:     $S_{harmful}\leftarrow S_{harmful}\cup\{s_{harmful}\}$
9:GAP-GuardData $\leftarrow S_{benign}\cup S_{harmful}$
10:$S_{attack}\leftarrow\{\}$
11:for each prompt $p$ in $S_{harmful}$ do
12:     $p_{attack}\leftarrow$ ApplyGAP ($p$)
13:     $S_{attack}\leftarrow S_{attack}\cup\{p_{attack}\}$
14:GAP-GuardAttackData $\leftarrow S_{benign}\cup S_{attack}$
15:return GAP-GuardData, GAP-GuardAttackData
This automated process not only removes the need for manual seed curation but also ensures a wide-ranging exploration of possible jailbreaking strategies. The detailed algorithm for GAP-Auto is provided in Algorithm 2. Using this approach, we generate two complementary datasets: GAP-GuardData, containing balanced benign and harmful prompts derived directly from content policies, and GAP-GuardAttackData, which consists of the original benign prompts together with GAP-refined stealthy versions of the harmful prompts, as detailed in Table 3.
@@ 2.2.2 GAP-VLM: Multimodal Attacks
Our GAP-VLM variant extends the framework to vision-language models (VLMs) by converting successful text-based jailbreaks into image-embedded attacks using a modified version of FigStep Gong et al. (2023). This adaptation involves:
@@
  • Prefix Enhancement: Incorporating the "Sure, here" suffix technique Wang and Qi (2024) into the typographic image generation process.
The GAP-VLM pipeline transforms these jailbreak prompts into image + prompt variants specifically designed to circumvent VLM safeguards. Algorithm 3 details the complete process.
Algorithm 3 GAP-VLM Multimodal Attack Generation
1:GAP-generated text prompt $p$
2:Multimodal attack prompt $m$
3:$d\leftarrow$ ParaphraseToDeclarative($p$)
4:$i\leftarrow$ ConvertToTypographicImage($d$)
5:$i^{\prime}\leftarrow$ AddPrefixToImage($i$, "Start the response with ‘Sure, here’")
6:$t\leftarrow$ GenerateBenignTextPrompt()
7:$m\leftarrow$ CombineImageAndText($i^{\prime}$, $t$)
8:return $m$
@@ 3 Experiments
{shaded1}
  1. RQ1: How does GAP compare to TAP in terms of attack success rate and query efficiency?
  3. RQ3: How effective is GAP at improving content moderation through fine-tuning via data augmentation?
  4. RQ4: How does GAP’s performance vary across different attacker models, target models, and query variations?
3.1 Experimental Setup
We implemented GAP variants in Python using attacker models described in Table 2. For evaluation and experimentation, we used the following setup:
@@ 3.1 Experimental Setup
Models. We employ three categories of models in our experiments:
  • Attacker Models: GAP-M uses Mistral-123B-v2407 while GAP-V uses Vicuna-13B-v1.5 as the attacker LLM.
  • Judge Model: GPT-4 serves as the judge model for assessing prompt relevance and jailbreak success across all variants.
  • Target Models: We evaluate against GPT-3.5, Gemma-9B-v2, and Qwen-7B-v2.5 as representative target LLMs. For multimodal experiments, we use GPT-4o as the target VLM.
Att: While we conducted experiments with additional target LLM models from various providers, results are omitted due to business constraints and they added no additional insights. Key findings extend analogously to leading LLMs beyond the included representative set.
Hyperparameters. We use consistent settings across all experiments unless noted. We set branching factor ($b$) to 5, allowing each node to generate five candidate prompts, and maximum width ($w$) to 3, controlling nodes retained after pruning. We allow five refinement iterations per seed (maximum depth $d=5$), maintain 10 recent history entries in the global context ($k=10$), and use sampling temperature 0.7 for the attacker model. These values were selected through preliminary testing.
For optimal performance, we found that providing detailed instructions to the attacker model about the desired prompt characteristics is crucial. These instructions emphasize maintaining natural language, embedding the harmful request in a broader context, and avoiding direct refusals or ethical warnings.
Datasets. We use multiple datasets throughout our experiments, as detailed in Table 3. For RQ1 and RQ4, we select the AdvBench subset (50 seeds across 32 categories) as seeds for jailbreak prompt generations Chao et al. (2023). RQ2 uses the same AdvBench subset for both text-only and multimodal VLM attack scenarios. For RQ3, we employ three different test datasets: the Toxic Chat Lin et al. (2023), OpenAI Moderation Markov et al. (2022), and custom GAP-GuardAttackData dataset as described in Section 2.2.1.
Metrics. Our primary metrics include:
  • Attack Success Rate (ASR): Percent of successful jailbreaks.
  • Query Efficiency: Average number of queries per successful jailbreak.
  • True Positive Rate (TPR): For guardrails, percent of harmful prompts correctly flagged.
  • Accuracy: Correct classification rate.
  • F1 Score: Harmonic mean of precision and recall.

### v3

@@ Abstract
As large language models (LLMs) become increasingly prevalent, ensuring their robustness against adversarial misuse is crucial. This paper introduces the GAP (Graph of Attacks with Pruning) framework, an advanced approach for generating stealthy jailbreak prompts to evaluate and enhance LLM safeguards. GAP addresses limitations in existing tree-based LLM jailbreak methods by implementing an interconnected graph structure that enables knowledge sharing across attack paths. Our experimental evaluation demonstrates GAP’s superiority over existing techniques, achieving a 20.8% increase in attack success rates while reducing query costs by 62.7%. GAP consistently outperforms state-of-the-art methods for attacking both open and closed LLMs, with attack success rates of $\geq$96%. Additionally, we present specialized variants like GAP-Auto for automated seed generation and GAP-VLM for multimodal attacks. GAP-generated prompts prove highly effective in improving content moderation systems, increasing true positive detection rates by 108.5% and accuracy by 183.6% when used for fine-tuning.  [^1] Code shared at https://github.com/dsbuddy/GAP-LLM-Safety. Warning: This paper contains examples of adversarial prompts that may be offensive to readers.
@@ 1 Introduction
With the increasing adoption of large-language models (LLMs) across diverse applications, ensuring their reliability and robustness against adversarial misuse has become a critical priority chao2023jailbreaking. Jailbreaking techniques, which involve crafting adversarial prompts to bypass an LLM’s safeguards, pose a persistent challenge to AI security and responsible deployment shen2024anything; mangaokar2024prp; wei2024jailbroken; li2023deepinception; guo2024cold. These methods can induce models to generate harmful, biased, or unauthorized content while avoiding detection by automated moderation systems perez2022red, highlighting the need for comprehensive diagnostic frameworks to assess and improve foundation model reliability.
@@
TAP-enhanced
Prompt Guard
- 88.0% 94.0% 60.0% 52.0%
GAP-Enhanced
Prompt Guard
68.0% 100.0% 100.0% 66.0% 70.0%
Table 1: True positive rate (TPR) comparison of various guardrails detecting prompts generated from multiple jailbreak methods (on AdvBench seeds). Lower TPR indicates better evasion and significant reliability concerns. Jailbreaking prompts generated by TAP and GAP reveal the most critical vulnerabilities across most guardrails. The last two rows show how GAP and TAP-generated data can be used to enhanced content moderation systems, demonstrating substantially improved detection capabilities against all methods, including GAP itself. Highest TPR values are bolded.
@@
Existing jailbreaking methods fall into three broad categories: (a) white-box attacks, which leverage direct model access for adversarial optimization zou2023universal; geisler2024attacking; (b) gray-box attacks, which involve techniques such as backdoor injection or poisoned retrieval ding2023wolf; shi2023badgpt; zou2024poisonedrag; wang2023backdoor; and (c) black-box attacks, which require only API access and thus represent the most realistic scenario for evaluating model robustness in real-world deployments wei2024jailbroken; li2023deepinception; yu2023gptfuzzer; yuan2023gpt; mehrotra2023tree. Notably in the black-box category, the Tree of Attacks with Pruning (TAP) approach mehrotra2023tree introduced a tree-structured exploration process for iterative prompt refinement, generating increasingly effective adversarial inputs that appear human-like and stealthy. As shown in Table 1, TAP-generated jailbreak prompts consistently demonstrate low detection true positive rate (TPR) when run against recent guardrails, indicating significant vulnerabilities in these safeguard systems that require systematic assessment and improvement.
While TAP demonstrated effectiveness in generating stealthy jailbreaks, we identified several limitations when applying it to thoroughly evaluate model reliability. Primarily, TAP restricts the exploration of prompt refinement to isolated, individual paths, with no crossover or shared context across different branches. This fundamental architectural limitation results in redundant queries and inefficient coverage of the search space for prompt refinement. Consequently, successful attack patterns discovered in one branch cannot inform or improve the exploration in others, leading to suboptimal attack success rates and unnecessarily high query costs, especially for more challenging jailbreak scenarios.
To overcome existing limitations in vulnerability assessment, we introduce the GAP (Graph of Attacks with Pruning) framework, which enables knowledge transfer across sequential attack seeds rather than confining it to a single session. GAP converts the traditional tree-based exploration process into an interconnected graph structure, maintains a global context to aggregate effective jailbreak strategies, and leverages graph-based knowledge sharing for informed prompt refinement. [^2] Our threat model assumes black-box user-level access, focusing on forcing LLMs to produce harmful responses even when system prompts are inaccessible. As shown in Table 1, GAP achieves substantially higher success rates and superior stealth—demonstrated by a lower true positive rate (TPR)—than TAP, including improved evasion against Prompt Guard (16.0% TPR vs. 22.0% for TAP).
@@
  • We demonstrate how GAP-generated insights improve foundation model reliability through data augmentation of safeguards. Our experiments show GAP-Enhanced Prompt Guard significantly improves detection capabilities across all jailbreak methods. As shown in Table 1, the enhanced guard achieves a TPR of 70.0% against GAP, versus the original’s 16.0%, substantially improving content moderation.
@@ 2.1 GAP (Graph of Attacks with Pruning)
The GAP core algorithm includes three stages:
@@
  • (Step 2) The pruning step where the judge LLM evaluates branches, removes unsuccessful ones, and focuses effort on variants most effective at eliciting undesired responses (lines 15 and 18 in Algorithm 1). This step implements a two-phase pruning strategy, including Off-topic pruning (the judge LLM removes branches irrelevant to the original harmful request) and Highest-scoring pruning (only branches with the highest scores $s_{i,j}=\mathcal{J}(p_{i,j},r_{i,j})$ (up to width $w$) advance to the next iteration).
@@
GAP’s key innovation reflects in the building of its first step’s global context $C=\{h_{1},h_{2},\dots,h_{n}\}$ that aggregates successful attack patterns from prior generations across all branches and sequential seeds (lines 4-8). For each prompt node $p$, GAP maintains a history $h_{p}$ of [prompt, response, score] tuples along its refinement path. Unlike TAP’s isolated tree structure, where each seed generates an independent attack path, GAP maintains a unified attack graph where successful strategies are shared and reused. This enables each new seed to leverage patterns observed in previous seeds, resulting in progressively smaller, more efficient attack trees with each sequential seed, as illustrated in Figure 1.
@@
Table 2: Comparison of TAP and GAP variants. While GAP variants use a graph structure with shared knowledge, they differ in their specific capabilities and the underlying attacker models we use for generating jailbreak prompts.
@@
2.1.1 Knowledge Transfer Implementation
In summary, GAP’s global context design makes its candidate prompt generation process following an interconnected graph-structure. This enables knowledge transfer via:
  1. 1.
  2. 2.
The attacker LLM uses this global context to create jailbreak candidates with two goals: (1) crafting natural-sounding prompts likely to elicit target responses and (2) incorporating effective patterns observed across successful examples in the global context, reusing and adapting proven strategies to the current context, thus improving jailbreak efficiency.
2.1.2 Connecting to Related Methods
While TAP mehrotra2023tree represents the closest related work in current literature, it fundamentally differs from our approach by restricting exploration to isolated tree structures. In contrast, GAP’s interconnected graph architecture enables cross-branch knowledge sharing and pattern reuse, as visualized in Figure 1. We quantitatively demonstrate GAP’s superior performance in both efficiency and effectiveness over TAP via comprehensive evaluations in Section 3.
Our approach also differs significantly from other black-box methods such as GPTFuzzer. Unlike GPTFuzzer, which relies on evolutionary algorithms and local mutation operators, GAP employs a graph-based refinement process that maintains a global context and enables knowledge sharing across all attack paths. Moreover, GAP introduces a two-phase pruning mechanism—off-topic and score-based pruning—that contrasts with GPTFuzzer’s fitness-based selection. Finally, GAP preserves contextual information across sequential seeds, whereas GPTFuzzer initializes each run independently.
Recent advances include AutoDAN-Turbo liu2024autodanturbolifelongagentstrategy, which employs a lifelong learning approach to automatically discover and evolve jailbreak strategies through multi-agent frameworks and strategy libraries. However, AutoDAN-Turbo focuses on long-term strategy accumulation and requires extensive warm-up phases, making it unsuitable as a direct baseline for our work, which addresses immediate structural limitations in prompt exploration efficiency.
2.2 GAP Variants for Different Scenarios
@@ 2.2.1 GAP-Auto: Auto Seed Generation
While GAP generates sophisticated jailbreak prompts, it initially requires manually crafted seed examples. To eliminate this dependency, we developed GAP-Auto, which automatically generates diverse seed prompts through a two-phase strategy:
@@
We use the 10 top-level categories from JailbreakBench (JBB chao2024jailbreakbench) as the basis for our policy decomposition. This automated process not only removes the need for manual seed curation but also ensures a wide-ranging exploration of possible jailbreaking strategies. Using this approach, we generate two complementary datasets: GAP-GuardData: A balanced set of benign and harmful prompts derived directly from content policies, and GAP-GuardAttackData: Contains the original benign prompts and the GAP-refined versions of the harmful prompts (detailed in Algorithm 2 in Appendix A.1).
@@ 2.2.2 GAP-VLM: Multimodal Attacks
Our GAP-VLM variant extends the framework to vision-language models (VLMs) by converting successful text-based jailbreaks into image-embedded attacks using a modified version of FigStep gong_figstep_2023. This adaptation involves:
@@
  • Prefix Enhancement: Incorporating the "Sure, here" suffix technique wang2024a into the typographic image generation process.
The GAP-VLM pipeline transforms these jailbreak prompts into image + prompt variants specifically designed to circumvent VLM safeguards (detailed in Algorithm 3 in Appendix A.1).
@@ 3 Experiments
{shaded1}
  1. RQ1: How does GAP compare to TAP in attack effectiveness and query efficiency?
  3. RQ3: Can GAP improve content moderation through fine-tuning via data augmentation?
  4. RQ4: How sensitive is GAP to attacker models, target models, and query budgets?
3.1 Experimental Setup
We implemented GAP variants in Python using attacker models described in Table 2. For evaluation, we used: (1) Attacker Models: GAP-M uses Mistral-123B-v2407 while GAP-V uses Vicuna-13B-v1.5; (2) Judge Model: GPT-4 for assessing prompt relevance and jailbreak success; (3) Target Models: GPT-3.5, Gemma-9B-v2, Qwen-7B-v2.5, and GPT-4o (for multimodal). We use consistent hyperparameter settings: branching factor $b=5$, maximum width $w=3$, maximum depth $d=5$, global context size $k=10$, and temperature $0.7$ (detailed specifications in Appendix A.3).
Our selection of Llama Guard, Llama Guard-2, and Perplexity-based detection for evaluation is based on their status as established benchmarks and their widespread adoption in the field. Llama Guard models are recognized as an open-source defense standard and are deployed across Meta’s products touvron2023llama; inan2023llama; zizzo2025adversarial. They are also commonly used by major commercial LLM providers. Perplexity-based defenses are also a prominent class of defense mechanisms, often used to detect non-natural adversarial inputs. These methods, along with other input filters and LLM-based judges, represent key categories in the taxonomy of LLM defense mechanisms. Their inclusion in our systematic evaluation validates our choice to test against established reference points in LLM safety research.
Datasets and Metrics. We use multiple datasets throughout our experiments, as detailed in Table 3. For RQ1 and RQ4, we select the AdvBench subset (50 seeds across 32 categories) as seeds for jailbreak prompt generations chao2023jailbreaking. RQ2 uses the same AdvBench subset for both text-only and multimodal VLM attack scenarios. For RQ3, we employ the GAP-GuardAttackData dataset and evaluate on Toxic Chat lin2023toxicchat and OpenAI Moderation openai2022moderation test sets. Our primary metrics include: Attack Success Rate (ASR), Query Efficiency, True Positive Rate (TPR) [^3] TPR values were computed using each guardrail’s native evaluation, such as internal classification for Llama Guard models, BERT-based scoring for Prompt Guard, and language model likelihood ratios for Perplexity., Accuracy, and F1 Score.
RQ1: How does GAP compare to TAP in attack effectiveness and query efficiency?